OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = -x + 1, passing through the point (4,1)?

OpenStudy (anonymous):

Okay, so it says we need to have a line parallel to the one above. Whenever lines are parallel that means they will have the same slope. :) That means we have y=-x so far and to get the rest of the equation we use the point-slope formula. y-y1=m(x-x1) In this problem m=slope, and (4,1) is (x1,y1). So once we pluck in the stuff in, we have- y-1=-1(x-4) Now we distribute the parenthesis y-1=-x+4 And now we add 1 to both sides :) y=x+5 -> This is the equation which we now found :)